the length of a rectangle is 2cm less than 3 times its width.if the perimeter of the rectangle is 68m find the dimensions of the rectangle?
Answer by mananth(16946)

let width be x
the length will 3x-2
2*x + 2(3x-2) = 68
2x+6x-4=68
8x=68+4
8x=72
x=9 m
length will be 3x-2 = 25 m